OpenStudy (anonymous):

Given that Sin A = 4/5, 0 < A < π/2 and Cos B = -12/13, π< B < 3π/2, find Cos 2A

OpenStudy (anonymous):

Cos 2A= 1-2 sin^2 A =1-2 (4/5)^2 =1-2(16/25) =1-32/25 =-7/25. I don't know why there is a need of Cos B.

OpenStudy (anonymous):

But Cos is positive in 0<A<90.So answer is 7/25

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no problem...